### v1

# Legilimens: Practical and Unified Content Moderation for Large Language Model Services

Abstract.
Given the societal impact of unsafe content generated by large language models (LLMs), ensuring that LLM services comply with safety standards is a crucial concern for LLM service providers. Common content moderation methods are limited by an effectiveness-and-efficiency dilemma, where simple models are fragile while sophisticated models consume excessive computational resources. In this paper, we reveal for the first time that effective and efficient content moderation can be achieved by extracting conceptual features from chat-oriented LLMs, despite their initial fine-tuning for conversation rather than content moderation. We propose a practical and unified content moderation framework for LLM services, named Legilimens, which features both effectiveness and efficiency. Our red-team model-based data augmentation enhances the robustness of Legilimens against state-of-the-art jailbreaking. Additionally, we develop a framework to theoretically analyze the cost-effectiveness of Legilimens compared to other methods.
We have conducted extensive experiments on five host LLMs, seventeen datasets, and nine jailbreaking methods to verify the effectiveness, efficiency, and robustness of Legilimens against normal and adaptive adversaries. A comparison of Legilimens with both commercial and academic baselines demonstrates the superior performance of Legilimens. Furthermore, we confirm that Legilimens can be applied to few-shot scenarios and extended to multi-label classification tasks.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 31th ACM Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, Utahisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
1. Introduction
With the rise of Large Language Models (LLMs), there has been a significant increase in global user engagement with online LLM services. However, the impressive and unknown capacities of chat-oriented LLMs can be used to malicious ends. Although LLMs are currently deployed on a large scale without properly taking into account the harm they can cause to society, there is a growing concern of and calls for compliance of LLMs to safety standard. Since model alignment by removing undesired behaviors from models has been empirically (Wallace et al., 2019; Yu and Sagae, 2021; Xu et al., 2021; Li et al., 2023; Liu et al., 2023; Zhu et al., 2023; Yuan et al., 2023) and theoretically (Wolf et al., 2023) proved to have fundamental limitations, service providers are tirelessly pursuing an effective and efficient content moderation solution.
However, existing content moderation methods have been found to be dangerously brittle. Even commercial solutions like OpenAI’s can be easily bypassed (Li et al., 2023; Liu et al., 2023; Zhu et al., 2023; Yuan et al., 2023) due to the inadequate ability of simple classifiers (Wei et al., 2023). Some researchers explored using LLMs as more sophisticated classifiers (Wang et al., 2023; Huang and Sun, 2023; Ma et al., 2023; Cao et al., 2023). But LLM-based content moderation imposes too great a burden on already insufficient computational resources (Griffin, 2023), rendering it impractical in real-world applications. This dilemma of effectiveness and efficiency requires urgent resolution.
Figure 1. Legilimens is a unified content moderation framework for almost all LLM services, which features both effectiveness and efficiency. By monitoring conceptual features generated in the regular inference process of LLMs, Legilimens achieves lightweight moderation.
Table 1. Legilimens versus existing works.
Method Model Architecture Model Parameters Extra Overhead* Domain
OpenAI Moderation API (Markov et al., 2023) GPT 117Msimilar-$\sim$1.5B $O(n)$ General content
[…]
BeaverDam-7B (Ji et al., 2023) LLaMA-7B 7B $O(n)$ QA
LLaMA Guard2 (Inan et al., 2023) LLaMA 3-8B 8B $O(n)$ QA
GradSafe (Xie et al., 2024) -§ -§ $O(n)$ General content
Wang et al(Wang et al., 2023) GPT-3 175B $O(n)$ Hate speech
Cao et al(Cao et al., 2023) GPT-4, LLaMA2 100T, 13B $O(n)$ General content
FakeGPT (Huang and Sun, 2023) GPT-3.5 $\geq$175B $O(n)$ Fake news
Ma et al(Ma et al., 2023) ChatGLM2-6B, Baichuan-13B 6B, 13B $O(n)$ General content
Legilimens (Ours) 1$\sim$3-Layer MLP 8$\sim$4.7M $\mathbf{O(1)}$ General content, QA
  • (i) †: The exact GPT architecture has not been disclosed (Markov et al., 2023). Based on the publication year (Yenduri et al., 2023) and the API response time, we speculate that it is either GPT-1 or GPT-2. (ii) ‡: The CNNs are distilled from trained BERT-based models. The exact architecture has not been disclosed. (iii) *: $n$ denotes the (token) length of input. The complexity is estimated based on the model architecture and our experimental results. (iv) ¶: “QA” refers to moderating question-answering (QA) data to assess the safety of the answers. (v) §: GradSafe calculates the difference in gradients between safe and unsafe prompts for detection, thus there is no detection model.
In this paper, we make the first attempt to balance effectiveness and efficiency in content moderation for chat-oriented LLMs. As shown in Figure 1, we propose a practical and unified content moderation framework for LLM services, named Legilimens  [^1] A legilimens is a person skilled in magically navigating through the many layers of a person’s mind and correctly interpreting one’s findings in the fantasy novels Harry Potter by J.K. Rowling (Rowling and Lauer, 2001), which extracts distinctive conceptual features from LLMs in a lightweight manner. Legilimens is built on the idea of leveraging the regular inference process of an LLM (the host LLM) to extract effective features for content moderation while incurring minimal overhead. A comparison of Legilimens with existing works is shown in Table 1. However, despite the simplicity and straightforwardness of the basic idea, realizing this concept into a practical system poses several challenges.
  • How to design an efficient and unified framework for input and output moderation that is universally applicable to almost all kinds of LLMs?
Our basic idea is to use the host LLM as a feature extractor. However, it is challenging to design a unified framework for LLMs of different architectures, for textual input or output of different lengths, and for different moderation tasks. First, LLMs can be categorized into Encoder-decoder and decoder-only by their architectures. And decoder-only LLMs can be further divided into causal decoders and prefix decoders. In addition, LLMs commonly have different model parameters, e.g., LLaMA2 has three sizes, 7B, 13B and 70B (Touvron et al., 2023b). Thus, the proposed method should be independent of model architectures and model sizes. Second, LLMs are sequence-to-sequence models that input and output variable-length texts. In this process, variable-length features in LLMs are generated. The proposed method should be able to deal with variable-length input/output and at best reduce the complexity to be independent of the input/output length, i.e., $O(1)$. Third, content moderation for LLM services consists of different tasks, including moderating the input based on the input, moderating the output based on the input and output, and moderating the input based on the input and output. The proposed method should be applicable to these tasks.
To address this problem, we design a decoder-based concept probing method that leverages the features generated in the decoding process of the first and the last output tokens. This decoding process is common to different LLM architectures and different input/output. By the attention mechanism, the features generated during the decoding of the first output token correspond to concepts related to the entire input. Similarly, the features generated during the decoding of the last output token correspond to concepts related to both the input and output. We utilize these two types of features to achieve both input and output moderation, namely $I$-moderation and $O$-moderation. Note that, our method focuses solely on two tokens, regardless of the input or output length. Thus, the complexity of our method is reduced to a constant level. We further demonstrate that Legilimens is applicable in few-shot scenarios, thereby reducing the cost for service providers to set up our content moderation system. In this way, we achieve an efficient and unified content moderation framework for LLM services.
  • How to defend against recent jailbreak attacks that aim to bypass the safety mechanisms of LLMs?
Jailbreaking refers to the process of circumventing the safety mechanisms placed on LLM services such as model alignment and content moderation. For instance, a common method to jailbreak model alignment is to instruct LLMs to emulate a “Do Anything Now” (DAN) behavior (King, 2023). An instance of jailbreaking content moderation involves performing orthographic transformations, such as Base64 encoding (Rao et al., 2024), on textual input or output to bypass content filters, as state-of-the-art LLMs can inherently process such encoded text. While LLM service providers have implemented stricter rules to prevent the use of such jailbreak prompts (Markov et al., 2023), they cannot entirely eliminate jailbreaking due to the numerous ways to construct prompts conveying the same meaning, owing to the inherent flexibility of natural languages.
To cope with this problem, we equip a red-team model-based data augmentation method in the training stage of Legilimens. Specifically, we employ a local LLM to work as a red-teaming model, generating highly adversarial jailbreak prompts from initially naive unsafe prompts. Through evaluating Legilimens against static and dynamic jailbreaking methods, we validate that incorporating augmented data during the training phase renders highly robust detection of unsafe content.
We have implemented a fully-functional prototype of Legilimens on five different host LLMs and evaluated its performance through extensive experiments on seventeen diverse datasets across three content moderation tasks (i.e., $I$-, $O$-, $IO$-moderation). Note that Although previous studies (Azaria and Mitchell, 2023; Chen et al., 2024) have utilized the hidden states of LLMs for hallucination detection, the potential of these hidden states for diverse content moderation tasks remains largely unexplored. Our work not only demonstrates that effective and robust content moderation can be achieved by extracting conceptual features from chat-oriented LLMs and augmenting prompts through red-teaming, but also underscores this paradigm’s ability to successfully navigate the trade-off between effectiveness and efficiency.
Our experiments, which include assessments against both static and dynamic jailbreaking, confirm the robustness of Legilimens against adversarially-designed prompts. Furthermore, the comparison with seven commercial and academic baselines demonstrates that Legilimens achieves the best performance in detecting unsafe content while maintaining the highest efficiency. Additionally, we validate that Legilimens can be applied to few-shot scenarios, and extended to perform multi-label classification tasks. We have open-sourced our code [^2] https://github.com/lin000001/Legilimens in a hope to incentivize more research in this area.
We summarize our contributions as follows:
  • We propose a practical and unified content moderation framework tailored for LLM services, which features a balance between effectiveness and efficiency.
  • We develop a concept probing technique that applies to mainstream encoder-decoder and decoder-only LLMs. Additionally, our red-team model-based data augmentation method enhances Legilimens to robustly resist jailbreaking.
  • We conduct extensive experiments to verify the effectiveness and efficiency of our content moderation framework against both non-adversarial and adversarial queries, outperforming seven commercial and academic baselines.
2. Background
2.1. Large Language Model
Typically, large language models (LLMs) refer to language models that contain billions of parameters, which are trained on massive text data (Zhao et al., 2023). Famous examples include GPT-3 (Brown et al., 2020a; Ouyang et al., 2022), GPT-4 (OpenAI, 2023a), GLM (Du et al., 2022; Zeng et al., 2023), LLaMA (Touvron et al., 2023a, b), etc. In this part, we introduce the basic components, architectures, and inference workflow of LLMs.
2.1.1. Basic Component
Given the excellent parallelizability and capacity, the Transformer architecture (Vaswani et al., 2017; Lin et al., 2021) has become the de facto backbone to almost all LLMs, making it possible to scale language models to hundreds or thousands of billions of parameters. The vanilla Transformer (Vaswani et al., 2017) is a sequence-to-sequence model and consists of an encoder and a decoder, each of which is a stack of $K$ identical blocks. Each encoder block is mainly composed of a multi-head self-attention module and a position-wise feed-forward network (FFN). For building a deeper model, a residual connection (He et al., 2016) is employed around each module, followed by layer normalization (Ba et al., 2016) module. Compared to the encoder blocks, decoder blocks additionally insert cross-attention modules between the multi-head self-attention modules and the position-wise FFNs. Furthermore, the self-attention modules in the decoder are adapted to prevent each position from attending to subsequent positions in the training phase.
2.1.2. Architecture
In general, the mainstream architectures of existing LLMs can be roughly categorized into two major types, namely encoder-decoder and decoder-only.
Encoder-Decoder Architecture. The vanilla Transformer model is built on the encoder-decoder architecture (Vaswani et al., 2017), which consists of two stacks of Transformer blocks as the encoder and decoder, respectively. The encoder encodes the input sequence for generating its latent representations, while the decoder performs cross-attention on these representations and generates the target sequence in an auto-regressive manner. So far, there are only a small number of LLMs that are built based on the encoder-decoder architecture, e.g., Flan-T5 (Chung et al., 2022).
Decoder-Only Architecture. Models of this type only have the decoder but no encoder. According to the self-attention mechanism used in the decoder, decoder-only models can be further divided into the causal decoder architecture and the prefix decoder architecture.
The causal decoder architecture incorporates the unidirectional attention mask, to guarantee that each input token can only attend to the past tokens and itself. The input and output tokens are processed in the same fashion through the decoder. As representative language models of this architecture, the GPT series models (Brown et al., 2020a; Ouyang et al., 2022; OpenAI, 2023a) are developed based on the causal decoder architecture. So far, the causal decoders have been widely adopted as the architecture of LLMs by various existing LLMs, such as LLaMA (Touvron et al., 2023a, b), Dolly (Conover et al., 2023b, a), and Falcon (Penedo et al., 2023).
The prefix decoder architecture (a.k.a., non-causal decoder) revises the masking mechanism of causal decoders, to enable performing bidirectional attention over the prefix tokens (Dong et al., 2019) and unidirectional attention only on generated tokens. In this way, like the encoder-decoder architecture, the prefix decoders can bidirectionally encode the prefix sequence and auto-regressively predict the output tokens one by one, where the same parameters are shared during encoding and decoding. Existing representative LLMs based on prefix decoders include GLM (Du et al., 2022; Zeng et al., 2023) and U-PaLM (Tay et al., 2023).
2.1.3. Model Inference
In typical model inference process of LLMs, a prompt $\mathbf{p}=p_{1}p_{2}\cdots p_{n}$ is fed into LLMs to generate a response sequence $\mathbf{r}=r_{1}r_{2}\cdots r_{m}$ auto-regressively, i.e.,
(1) $r_{i+1}$ $=\mathop{\arg\max}\limits_{r}\mathbb{P}\left(r|\mathbf{p}\oplus\mathbf{r}_{1:i}\right)\quad\text{({greedy search})}$
$\text{{or}}\quad r_{i+1}$ $\sim\mathbb{P}\left(r|\mathbf{p}\oplus\mathbf{r}_{1:i}\right)\quad\text{({sampling-based methods})},$
where $\oplus$ denotes concatenating the previous output tokens to the end of the input sequence until a special sentence ending token (usually denoted as [eos]) is generated. The first decoding method is greedy search, which predicts the most likely token at each step based on the previously generated tokens. The other decoding method is sampling, which randomly samples the next token based on the probability distribution to enhance the randomness and diversity during generation.
From the inference process and the self-attention mechanism of LLMs we know that LLMs output the first token $r_{1}$ of $\mathbf{r}$ leveraging the information of $\mathbf{p}$, and output the last token (i.e., [eos]) with the information of both $\mathbf{p}$ and $\mathbf{r}$, i.e.,
(2) $r_{1}=\mathcal{H}(\mathbf{p})\quad\text{and}\quad\text{{[eos]}}=\mathcal{H}(\mathbf{p}\oplus\mathbf{r}),$
where $\mathcal{H}(\cdot)$ denotes the inference function of LLMs. We utilize this inference process as a feature extractor for the downstream content moderation task, which we elaborate on in §4.
2.2. Content Moderation
To mitigate potential harm and misuse of LLMs, two safety mechanisms are commonly applied, i.e., model alignment involves training LLMs to reject unsafe prompts, while content moderation employs filters to block unsafe prompts and responses.
2.2.1. Unsafe Content
Given that out-of-the-box LLMs have the potential to generate misinformation, propagate harmful content, or produce unintended responses with significant negative societal impact (Weidinger et al., 2022; Cui et al., 2024; Barrett et al., 2023; Weidinger et al., 2021), content moderation is essential for identifying such unsafe content within prompts and responses. The definitions of unsafe content in previous literature often include profanities, identity attacks, sleights, insults, threats, sexually explicit content, demeaning language, and language that incites violence (Schmidt and Wiegand, 2017; Fortuna and Nunes, 2018; Gorwa et al., 2020). In practice, different service providers may adopt different definitions and taxonomies of unsafe content.
2.2.2. Moderation Tasks
Content moderation for LLMs involves moderating the input prompt and the output response, referred to as $I$-moderation and $O$-moderation respectively. In $I$-moderation (resp. $O$-moderation), the moderator assesses whether the input prompt (resp. output response) is unsafe, based solely on the input prompt (resp. output response) or on the entirety of the prompt and response. The combined task, referred to as $IO$-moderation, aims to assess both the input prompt and the output response.
2.3. Jailbreak
Jailbreaking is a process that employs adversarially-designed prompts to circumvent the safety mechanisms imposed on LLMs by their service providers. Several efforts have been made to taxonomize jailbreaking (Liu et al., 2023; Rao et al., 2024; Shen et al., 2023; Chu et al., 2024), based on which jailbreaking can be categorized into semantic transformation and syntactic transformation.
  • Semantic transformation involves manipulating the semantics of prompts. Based on the patterns of semantic manipulation, three general types of semantic transformation have been identified (Liu et al., 2023).
    • Pretending: This type of prompts tries to alter the conversation background or context while maintaining the same intention, e.g., creating a role-playing game context.
    • Attention Shifting: This type of prompts aims to change both the conversation context and intention, e.g., shifting the intention of the prompt from asking the model questions to making it construct a paragraph of text.
    • Privilege Escalation: This type of prompts seeks to directly circumvent the imposed restrictions. For instance, “Do Anything Now” prompts mentioned in §1 belong to this category.
  • Syntactic transformation modifies only the syntax of prompts without altering the semantics. Examples include string splicing, common encoding, and simple encryption. This type of transformation can effectively bypass rule-based filters.
Note that the generation of jailbreak prompts can be static and dynamic. Static prompts do not leverage response information from the target LLM while dynamic methods use previous responses as a reference to adversarially refine the prompt for launching the next attack.
2.4. Design Goal
We first define the system model in terms of the adversary and the defender, and then elaborate the design goals of Legilimens under the defined system model.
Adversary. The adversary attempts to induce undesirable behavior of LLM services, whether unintentionally or intentionally. This may include generating inappropriate content, disclosing sensitive information, or performing actions against programming constraints. To achieve this, the adversary may optimize the semantics or syntax of their prompts to outsmart the safeguards of LLMs by any jailbreak methods. The adversary can observe the returned responses and leverage this knowledge to refine the prompt in order to achieve their attack objectives.
Defender. The defender aims to ensure both the safety and helpfulness of the LLM service. To achieve this, the defender strives to apply content moderation mechanisms that are as accurate as possible, with full access to the host LLM.
Given this system model, we delineate two major goals of Legilimens.
  • Effectiveness. Legilimens should accurately identify unsafe prompts given to the host LLM or responses generated by the host LLM, even when confronted with jailbreak prompts.
  • Efficiency. Legilimens should introduce minimal overhead compared to the original inference process of the host LLM. Ideally, the overhead of Legilimens should not increase with the length of prompts or responses, i.e., maintaining a constant complexity of $O(1)$.
Figure 2. Design of Legilimens. Legilimens utilizes the conceptual features during the decoding of the first token to accomplish $I$-moderation, and the conceptual features during the decoding of the last token to achieve $O$- and $IO$-moderation.
3. Problem Formulation
In this section, we first formulate content moderation as a classification problem. Then we materialize the classification problem for three distinct moderation tasks.
3.1. Classification Formulation
A textual input $\mathbf{x}=x_{1}x_{2}\cdots x_{n}$ can be tokenized into $n$ frames of token embeddings, i.e., $\mathbf{X}\in\mathbbm{R}^{n\times d}$, where $d$ denotes the dimension of each token embedding. A typical content moderation problem is to determine whether $\mathbf{x}$ is unsafe by a moderator $\mathcal{M}_{\theta}(\cdot):\mathbbm{R}^{n\times d}\rightarrow\mathbbm{Y}$, parametrized by $\theta$, which can be expressed as
(3) $\mathop{\min}\limits_{\theta}\mathop{\mathbb{E}}\limits_{\mathbf{X},y\sim\mathbbm{D}}\mathcal{L}\left(\mathcal{M}_{\theta}(\mathbf{X}),y\right),$
where $\mathbbm{D}$ is a labelled content moderation dataset. $y\in\mathbbm{Y}$ is the ground-truth label of the corresponding input $\mathbf{X}$, denoting whether $\mathbf{X}$ is unsafe. $\mathcal{L}$ represents the loss function, typically measured using cross-entropy loss, which evaluates the accuracy of $\mathcal{M}_{\theta}(\mathbf{X})$.
In the classification problem above, when the training dataset $\mathbbm{D}$ is of sufficient size, the accuracy of the moderator $\mathcal{M}_{\theta}(\cdot)$ generally increases with the scale of its parameters $\theta$, but at the cost of increased computational complexity. Thus, efficiency and effectiveness are often in a trade-off relationship.
In the scenario of LLM services, the host LLM is expected to infer input prompts and generate corresponding responses. Given the substantial capacity of the host LLM, we aim to utilize the original inference process of the host LLM to materialize a portion of $\mathcal{M}_{\theta}(\cdot)$. This approach allows representative features to be extracted in a “free-riding” manner, i.e.,
(4) $\mathcal{M}_{\theta}\left(\mathbf{X}\right)=\mathcal{C}_{\theta\backslash\psi}\circ\mathcal{H}_{\psi}\left(\mathbf{X}\right),$
where $\mathcal{H}_{\psi}$ represents a portion of the host model parameterized by $\psi$, and $\mathcal{C}_{\theta\backslash\psi}$ denotes an additional classifier parameterized by $(\theta\backslash\psi)$. In this way, the formulation in Equation (3) is transformed into
(5) $\mathop{\min}\limits_{\theta\backslash\psi}\ \mathop{\mathbb{E}}\limits_{\mathbf{X},y\sim\mathbbm{D}}\mathcal{L}\left(\mathcal{C}_{\theta\backslash\psi}\circ\mathcal{H}_{\psi}\left(\mathbf{X}\right),y\right).$
Notably, since $|\psi|$ is sufficiently large, $|\theta|$ can be large when $|\theta\backslash\psi|=|\theta|-|\psi|$ is relatively small. Consequently, this approach potentially achieves a win-win scenario in terms of efficiency and effectiveness.
3.2. Materialization Tasks
Based on the object to be moderated, we materialize Legilimens into three moderation tasks, as depicted in Figure 2.
  • I-Moderation. In $I$-moderation, the object to be moderated is solely input prompts to be fed into the host LLM. The task is to determine whether the input prompts are unsafe, denoted as $\hat{y}=\mathcal{M}_{\theta}\left(\mathbf{P}\mid*\right)$, where $\mathbf{P}$ represents the token embeddings of prompt $\mathbf{p}$, and * signifies any additional information available to the moderator, e.g., output responses.
  • O-Moderation. In $O$-moderation, the object to be moderated is solely output responses generated by the host LLM. The task is to determine whether the output responses are unsafe, denoted as $\hat{y}=\mathcal{M}_{\theta}\left(\mathbf{R}\mid*\right)$, where $\mathbf{R}$ represents the token embeddings of response $\mathbf{r}$, and * signifies any extra information available to the moderator, e.g., input prompts.
  • IO-Moderation. In $IO$-moderation, the object to be moderated encompasses both input prompts and output responses, which is a combined task of $I$- and $O$-moderation, i.e., $\hat{y}=\mathcal{M}_{\theta}\left(\mathbf{P},\mathbf{R}\mid*\right)$.
The three tasks above can be applied along with different moderation strategies. In the case of $I$-moderation, the LLM service can halt once the prompt is determined to be unsafe, thereby saving subsequent computation. However, this strategy may potentially impair the overall helpfulness of the LLM service. Conversely, for $O$-moderation, the service halts until the response is determined to be unsafe. This strategy preserves helpfulness when prompts are unsafe but successfully handled by the host LLM, providing safe responses. Nevertheless, this strategy may require additional computational resources. $IO$-moderation halts the service upon detecting unsafe prompts or responses, representing a more rigorous moderation strategy compared to the first two tasks.
4. Legilimens: Design Details
In this section, we materialize $\mathcal{H}_{\psi}$ and $\mathcal{C}_{\theta\backslash\psi}$ in Equation (5) in the scenario of LLM services by concept probing and lightweight moderator construction.
4.1. Concept Probing
4.1.1. Conceptual Feature Extraction
Typical LLMs are constructed by stacking $K$ encoder-decoder or decoder-only Transformer blocks together, i.e.,
(6) $\mathcal{H}=\mathcal{H}_{K}\circ\mathcal{H}_{K-1}\circ\cdots\circ\mathcal{H}_{1},$
where $\mathcal{H}_{i}$ denotes the $i$-th block of the host LLM. To decode a token, the original inference process derives side features as follows,
(7) ~$\widetilde{\mathcal{H}}_{k}(\mathbf{X})\triangleq\left(\mathcal{H}_{k}\circ\mathcal{H}_{k-1}\circ\cdots\mathcal{H}_{1}\right)(\mathbf{X}),\quad 1\leq k\leq K.$
Our intuition is that LLMs develop concepts of the input along the inference process. Thus, we attain comprehensive conceptual features of $\mathbf{X}$ by fusing/concatenating the side features derived from the last several blocks, i.e.,
(8) ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{X})\triangleq\left[\widetilde{\mathcal{H}}_{(K-m)}(\mathbf{X})\oplus\cdots\oplus\widetilde{\mathcal{H}}_{K}(\mathbf{X})\right].$
The reason for deriving conceptual features from the last several blocks is to utilize more capacity of the host LLM, given that ~$\widetilde{\mathcal{H}}_{(K-m):K}$ can be viewed as containing all parameters of the host LLM. Note that, no additional computation is required to obtain the side features in Equation (8).
4.1.2. Probing Position
During the inference process, LLMs generate responses in an auto-regressive manner as mentioned in §2.1.3, where varied-length side features are derived. Managing these varied-length side features may lead to increased overhead as the response length grows. To reduce the complexity of Legilimens to $O(1)$, we resort to the attention span of self-attention mechanism.
As shown in Figure 2, the attention span during the decoding of the first token $r_{1}$ encompasses the entire prompt, indicating that the conceptual features derived in this process, ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P})$, encapsulate a summarized knowledge of prompt $\mathbf{p}$. Similarly, the attention span during the decoding the last token [eos] encompasses both the prompt and the response, encapsulating a summarized knowledge of prompt $\mathbf{p}$ and response $\mathbf{r}$ in ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P}\oplus\mathbf{R})$. Note that, both ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P})$ and ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P}\oplus\mathbf{R})$ are of a constant size of $m\times d_{\text{model}}$, where $d_{\text{model}}$ is the output dimension of Transformer blocks. Reasoning on these two conceptual features only introduces constant overhead.
4.2. Lightweight Moderator
4.2.1. Architecture
In this part, we construct a lightweight classifier to materialize $\mathcal{C}_{\theta\backslash\psi}$. Based on the comprehensive representation summarized by the host LLM, we find that a simple multi-layer perceptron (MLP) with few parameters is adequate for accurate content moderation, denoted as $\mathcal{C}_{\theta\backslash\psi}:\mathbbm{R}^{m\times d_{\text{model}}}\rightarrow\mathbbm{Y}$. Therefore, the entire moderator is composed in this way: $\mathcal{M}_{\theta}=\mathcal{C}_{\theta\backslash\psi}\circ\widetilde{\mathcal{H}}_{(K-m):K}$. We validate the effectiveness of our lightweight moderator in §5.2 compared with five baselines.
4.2.2. Model Training
Within the unified framework of concept probing, we train Legilimens to handle three content moderation tasks, i.e., $I$-, $O$-, $IO$-moderation, in accordance with Equation (5).
For $I$-moderation, we curate a training set of $(\mathbf{P},y_{p})$ to train the moderator, where the ground-truth label $y_{p}$ is assigned based on the safety of prompt $\mathbf{p}$. For $O$-moderation, as the original inference process of the host LLM does not involve reasoning on $\mathbf{R}$, we prepare a training set of $(\mathbf{P}\oplus\mathbf{R},y_{r})$ instead, where the ground-truth label $y_{r}$ is assigned based on the safety of response $\mathbf{r}$. Similarly, for $IO$-moderation, we prepare a training set of $(\mathbf{P}\oplus\mathbf{R},y_{p}|y_{r})$ to train the moderator.
4.3. Model-Based Data Augmentation
An adaptive adversary may employ jailbreaking to compromise the protection of Legilimens. Given that jailbreaking can alter the conceptual features of prompts and responses in order to evade detection, we augment the training data of Legilimens to bolster its resistance against jailbreaking. Specifically, we employ LLaMa2 as a red-teaming model $\mathcal{T}$, prompting it to rewrite naive unsafe prompts $\mathbf{p}$ into adversarially-designed jailbreak prompts $\mathbf{p}^{\prime}=\mathcal{T}(\mathbf{p})$. Our system prompt consists of three segments. The first segment is ”Do Anything Now”, aimed at directing the large model $\mathcal{T}$ to violate its principle. The second segment specifies the output format and rewriting instructions, derived from extensive research (e.g., (Liu et al., 2023; Rao et al., 2024; Xu et al., 2024; Chu et al., 2024)), systematically enumerating various jailbreaking methodologies, such as ”Character Role Play”, ”Logical Reasoning”, among others. The final segment provides a few examples through a few-shot demonstration while utilizing the ability of in-context learning (Brown et al., 2020b). We rewrite 20% of the unsafe prompts in the training set to enhance the robustness of Legilimens. In §5.4, we validate the robustness of Legilimens against state-of-the-art static and dynamic jailbreak attacks.
5. Evaluation
5.1. Setup
5.1.1. Prototype
We have implemented a prototype of Legilimens on the PyTorch (Paszke et al., 2019) platform and train the moderators according to Equation (5) using a single NVIDIA 3090 GPU. We set the default Legilimens configurations as $m=1$, and $\mathcal{C}_{\theta\backslash\psi}$ is a three-layer MLP. In the training phase, we utilize an Adam (Kingma and Ba, 2015) optimizer to update the parameters of the MLP-based classifier for 50 epochs with a learning rate of 1e-4, a weight decay ($\ell_{2}$ penalty) rate of 1e-3, and a batch size of 256.
5.1.2. Host Model
We apply Legilimens to five host LLMs with various architectures, i.e., ChatGLM3-6B (Zeng et al., 2023), LLaMA2-7B (Touvron et al., 2023b), Falcon-7B (Penedo et al., 2023), Dolly-7B-v2 (Conover et al., 2023b), Vicuna-7B-v1.5 (Zheng et al., 2023).
  • ChatGLM3. ChatGLM3 (Zeng et al., 2023) is an open-source bilingual (English and Chinese) LLM following a prefix decoder architecture, which utilizes a multi-query attention mechanism (Shazeer, 2019) and the SwiGLU (Shazeer, 2020) activation function. ChatGLM3 comprises 28 Transformer blocks.
  • LLaMA2. LLaMA2 (Touvron et al., 2023b) is an open-source LLM developed by Meta using a causal decoder architecture, employing a grouped-query attention and the SwiGLU activation function. LLaMA2 consists of 32 Transformer blocks.
  • Falcon. Falcon (Penedo et al., 2023) is an open-source LLM using a causal decoder architecture, incorporating a multi-query attention mechanism and the GeLU (Dauphin et al., 2017) activation function. Falcon comprises 32 Transformer blocks.
  • Dolly. Dolly (Conover et al., 2023b) is fine-tuned from EleutherAI’s Pythia-6.9B (Biderman et al., 2023) on an instruction-tuning dataset of around 15,000 samples. It uses a causal decoder architecture, sparse attention mechanism (Child et al., 2019) and the GeLU activation function. Dolly comprises 32 Transformer blocks.
  • Vicuna. Vicuna (Zheng et al., 2023) is fine-tuned from LLaMA2 on around 125,000 instruction-tuning samples. The architecture of Vicuna is the same as LLaMA2.
The output dimension $d_{\text{model}}$ of Falcon is 4544, while it is 4096 for the other four LLMs.
5.1.3. Baselines
We utilize four commercial and three academic state-of-the-art content moderation methods as our baselines, i.e., Google Perspective API (Jigsaw, 2017) (commercial), OpenAI Moderation API (OpenAI, 2023b) (commercial), BeaverDam (Ji et al., 2023) (academic), LLaMA Gurad2 (Inan et al., 2023) (academic), GradSafe (Xie et al., 2024) (academic), GPT-3.5-Turbo (Brown et al., 2020a) (commercial), and GPT-4 (OpenAI, 2023a) (commercial). Detailed information about the first five baselines is listed in Table 1. To adapt GPT-3.5-Turbo and GPT-4 for content moderation, we employ manually designed prompts as illustrated in Appendix B. We will demonstrate that Legilimens outperforms these seven baselines in §5.2.
5.1.4. Metrics
Two typical and standard metrics are used to evaluate the content moderation performance of Legilimens.
  • Accuracy (ACC) is a measure of the correctness of content moderation, calculated as the ratio of the number of correct predictions to the total number of predictions made.
  • Area Under the ROC Curve (AUC) measures the area underneath the ROC (Receiver Operating Characteristic) curve. The AUC value ranges from 0 to 1, where a higher value indicates better capability to distinguish between safe and unsafe content.
  • False Positive Rate (FPR) is a measure of the ratio of falsely predicted positive instances to the total actual negative instances. FPR signifies the proportion of safe content that is incorrectly classified as unsafe in content moderation.
  • False Negative Rate (FNR) represents the ratio of falsely predicted negative instances to the total actual positive instances. In the realm of content moderation, FNR denotes the proportion of unsafe content that is erroneously classified as safe.
5.1.5. Datasets
Our experiments involve seventeen datasets covering various tasks across different domains, as shown in Table 24 in Appendix G. Among these, four datasets are dedicated to traditional content moderation, while the others are specialized for LLM services.
  • HateXplain (Mathew et al., 2021): This dataset serves as a benchmark for hate speech detection, sourced from Twitter and Gab. It has been annotated by Amazon Mechanical Turk (MTurk) workers, who assigned three labels, i.e., hate, offensive, or normal. In our paper, we consider hate and offensive posts as unsafe, and the normal posts as safe.
  • Measuring Hate Speech (Kennedy et al., 2020): This dataset comprises comments from social media platforms. These comments have been labelled by MTurk workers, and the labels have been converted into a continuous score. We consider comments with a score over 0.5 as unsafe, and those with a score less than -1 as safe.
  • OIG-Safety (Nguyen et al., 2023): We utilize a subset of samples from OIG-Safety, labelled as casual and needs intervention, comprising 21,769 samples. We consider casual as safe and needs intervention as unsafe.
  • Jigsaw (Google Jigsaw, 2017): This dataset comprises seven categories of samples, namely innocent, severe toxicity, obscene, identity attack, insult, threat, and sexual explicit. We consider innocent as safe and the remaining categories as unsafe.
We also assess Legilimens using datasets specifically designed for $O$-moderation in LLM scenarios.
  • BeaverTails (Ji et al., 2023): We conduct a voting process on the original dataset to obtain our dataset, which consists of 110,822 question-answer pairs across 14 potential harm categories. Note that the prompts are all unsafe. We utilize this dataset for $O$-moderation.
  • BeaverTails-adv: This dataset is created by inserting the original prompts from BeaverTails into four types of jailbreak templates, including pretending, attention shifting, privilege escalation, and syntactic transformation. The jailbreak templates are collected from the Internet.
Table 2. Accuracy performance on various ${}^{\text{\textdagger}}$ and ${}^{\text{\textdaggerdbl}}$ compared with baselines.
Method I-Moderation (ACC, %) O-Mod. (ACC, %) IO-Mod. (ACC, %) Time/Query (ms)
HAT MHS OIG JIG BEA BEA-adv BAG BAG-adv
OpenAI Moderation 70.85 72.30 65.85 76.70 51.70 47.95 53.90 53.4 566.951
Perspective API 64.79 75.1 60.56 78.26 47.72 45.95 51.78 48.50 90.336
BeaverDam-7B 66.65 74.15 63.75 67.50 89.50 73.75 74.95 61.75 30.121
LLaMA Guard2 71.40 77.85 76.00 50.95 77.38 75.70 68.40 66.15 430.923
GradSafe 69.82 67.70 78.90 66.00 73.80 57.10 75.50 76.90 395.212
GPT-3.5-Turbo 68.00 75.52 60.70 52.50 65.73 55.99 55.82 44.42 681.800
GPT-4 74.00 75.00 79.35 68.00 83.00 75.00 73.00 60.26 801.720
[…]
  • †: Task alias: $I$-Moderation ($I$-Mod.), $O$-Moderation ($O$-Mod.) and $IO$-Moderation ($IO$-Mod.).
  • ‡: Dataset alias: HateXplain (HAT), Measuring Hate Speech (MHS), OIG-Safety (OIG), Jigsaw (JIG), BeaverTails (BEA), BeaverTail-adv (BEA-adv), BEA&AG (BAG), BEA-adv&AG (BAG-adv).
We create two datasets for $IO$-moderation in LLM scenarios since no relevant datasets are readily available.
  • BEA&AG: We compile this dataset by merging BeaverTails with an instruction-tuning dataset called Alpaca-GPT4 (Peng et al., 2023). As all prompts from BeaverTails are deemed unsafe, an equal number of samples from Alpaca-GPT4, considered safe, have been incorporated.
  • BEA-adv&AG: We compile this dataset by merging BeaverTails-adv with an equal number of samples from Alpaca-GPT4.
We also utilize various unseen datasets to assess the performance of Legilimens in scenarios involving potential distribution shifts.
  • BEA&PIQA: We combine non-repetitive prompts from BeaverTails and an equal number of prompts from the PIQA (Bisk et al., 2020) dataset. The latter are considered safe.
  • HarmBench (Mazeika et al., 2024): This dataset contains 320 human-written unsafe instructions, covering 7 semantic categories of behavior: cybercrime & unauthorized intrusion, chemical & biological weapons/drugs, copyright violations, misinformation & disinformation, harassment & bullying, illegal activities, and general harm.
  • SimpleSafetyTests (Vidgen et al., 2023): This dataset comprises 100 human-written unsafe simple questions or instructions, covering 5 harm areas: (1) suicide, self-harm, and eating disorders, (2) physical harm, (3) illegal and highly regulated items, (4) scams and fraud, and (5) child abuse.
  • MaliciousInstructions (Bianchi et al., 2023): This dataset contains 100 machine-written instructions generated by GPT-3 (text-davinci-003).
  • JADE (Zhang et al., 2023): The dataset contains 80 machine-written unsafe prompts, which were created by linguistic fuzzing to generate challenging prompts for evaluating LLM safety.
  • HExPHI (Qi et al., 2023): This dataset is sampled from AdvBench (Zou et al., 2023) and AnthropicRedTeam (Ganguli et al., 2022) and then refined manually and with LLMs. There are 330 unsafe instructions.
  • TDCRedTeaming (Mazeika et al., 2024): This dataset contains 50 human-written red-teaming instructions, covering 7 categories: bigotry & abusive language, violent content & conduct, illegal activities, malware & exploits, scams, misinformation & disinformation, other undesirable content.
We further employ two datasets as input of dynamic jailbreaking to stress test Legilimens in §5.4.2.
  • AdvBench: This dataset contains 50 prompts designed to elicit harmful information across 32 categories (Chao et al., 2023).
  • AdvBEA: This dataset contains 30 prompts sourced from the BeaverTails test set, encompassing categories such as violence, privacy, weapons, child abuse, and more.
5.2. Overall Performance
5.2.1. Effectiveness
In this part, we evaluate the overall effectiveness of Legilimens on eight datasets and three moderation tasks in comparison with seven baselines. We train Legilimens (hosted in LLaMA2) on the training set of each dataset and evaluate its performance on the corresponding test set. The results are presented in Table 2, Table 11 and Table 12, with the latter two in Appendix C.
For $I$-moderation, we utilize four datasets, i.e., HateXplain, MHS, OIG-Safety, and Jigsaw. As shown in Table 2, Legilimens achieves the best performance on all four datasets, including two for hateful speech and two for general unsafe content, with an accuracy of 82.19%, 90.48%, 94.67%, and 88.39%. Legilimens outperforms the second place by 8.19$\sim$15.32%. It validates the effectiveness of Legilimens on $I$-moderation across different domains.
For $O$-moderation, we assess eight methods on BeaverTails (normal version) and BeaverTails-adv (jailbreak version). Referring to Table 2, Legilimens achieves the second best performance on BeaverTails and the best performance on BeaverTails-adv. It is noteworthy that BeaverDam-7B has been trained end-to-end on the training set of BeaverTails, whereas only a small portion of Legilimens (the lightweight classifier) has been trained. Despite this, the performance gap between Legilimens and BeaverDam-7B is merely 1.39%. In contrast, OpenAI Moderation API, Perspective API, and GPT-3.5-Turbo perform roughly as well as random guessing. One possible reason is that these three baselines can not handle $O$-moderation when prompts are already unsafe. In addition, when we modify the unsafe prompts into jailbreak ones, the performance of BeaverDam-7B drops 15.75% while Legilimens only experiences a drop of 3.62%. This suggests that jailbreak templates have a negative influence on BeaverDam-7B in determining the safety of responses. In comparison, Legilimens demonstrates robustness against jailbreak prompts.
For $IO$-moderation, we evaluate eight methods on BEA&AG and BEA-adv&AG datasets. As depicted in Table 2, Legilimens achieves the best performance on both datasets with an accuracy of 99.56% and 97.34%, significantly outperforming the second place by 24.06% and 20.44%. OpenAI Moderation API, Perspective API, and GPT-3.5-Turbo still exhibit unsatisfactory performance. This suggests that these three methods may not be suitable for content moderation in LLM scenarios where both prompts and responses should be taken into account.
Table 3. Generalization performance ($I$-Moderation).
Test Set Host Model (ACC, %)
ChatGLM3 LLaMA2 Falcon Dolly Vicuna
BPI 97.53 98.58 97.82 96.12 98.03
HAB 90.94 79.69 59.38 82.19 73.75
SST 99.00 100.00 95.00 96.00 99.00
MAI 99.00 100.00 87.00 97.00 99.00
JAD 90.00 88.75 83.75 91.25 83.75
HEP 98.79 96.06 80.91 91.21 91.52
TDC 96.00 86.00 62.00 88.00 84.00
  • †: Legilimens is trained on the training set of BPI and test on the test sets of BPI and other unseen datasets.
  • ‡: Dataset alias: BEA&PIQA (BPI), HarmBench (HAB), SimpleSafetyTests (SST), MaliciousInstructions (MAI), JADE (JAD), HExPHI (HEP), TDCRedTeaming (TDC).
Table 4. Accuracy robustness against LLM-targeted static jailbreaking ($IO$-Moderation).
Jailbreaking Type ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
Pretending 99.529 0.9997 99.860 1.0000 99.557 0.9998 98.327 0.9978 99.506 0.9997
Attention Shifting 99.515 0.9997 99.860 1.0000 99.342 0.9994 97.954 0.9973 90.595 0.9961
Privilege Escalation 99.497 0.9997 99.846 1.0000 99.393 0.9996 94.901 0.9947 97.991 0.9986
Syntactic Transformation 93.298 0.9952 89.575 0.9971 97.338 0.9973 84.937 0.9776 87.221 0.9860
Overall 98.192 0.9987 97.339 0.9993 98.768 0.9987 93.983 0.9918 93.872 0.9953
Table 5. FPR and FNR robustness against LLM-targeted static jailbreaking ($IO$-Moderation).
Jailbreaking Type ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
Pretending 0.471 0.099 0.481 0.000 1.166 0.072 2.009 2.713 0.413 0.335
Attention Shifting 0.471 0.072 0.481 0.009 1.369 0.226 1.999 1.936 0.413 16.896
Privilege Escalation 0.471 0.072 0.481 0.000 1.225 0.217 1.846 3.654 0.413 10.013
Syntactic Transformation 0.471 15.304 0.481 13.296 1.263 5.048 1.999 29.134 0.413 22.395
Overall 0.461 4.016 0.346 3.763 1.206 1.583 1.933 9.235 0.413 12.229
5.2.2. Efficiency
We measure the additional overhead of moderating all samples in the test set of Jigsaw and present the average time per query for eight methods in Table 2. Legilimens only requires 0.003 milliseconds per query, which is over 10,040×10,040$\times$10 , 040 × faster than the other methods. This is because Legilimens only introduces the overhead of a three-layer MLP classifier, as most of the computation for Legilimens is completed in the original inference process of the host LLM.
To further examine the time complexity of the eight methods, we vary the length of inputs to be moderated and plot the corresponding computation times for OpenAI Moderation, Perspective API, BeaverDam-7B, and Legilimens in Figure 4 in Appendix F. The slopes of the fitted lines for the four methods are 2e-2, 3e-3, 4e-4, and -1e-8, respectively. Observing the trend of the computation times, we find that Legilimens exhibits a constant complexity of $O(1)$, while the computation times of the other methods increase as inputs lengthen. We do not plot the computation times of GPT-3.5-Turbo and GPT-4 because their complexity is assuredly not constant.
5.3. Generalization to Unseen Datasets
In this part, we evaluate the performance of Legilimens when it encounters potential distribution shifts. We train Legilimens on BEA&PIQA dataset, test it on six unseen datasets mentioned in §5.1.5, and present the results in Table 3. We observe that Legilimens, when hosted in ChatGLM3, LLaMA2, Dolly and Vicuna, generalizes well to most of the unseen datasets, achieving an accuracy of over 90.00%, 79.69%, 82.19%, 73.75% respectively. Although Legilimens in Falcon does not generalize well to HarmBench and TDCRedTeaming, it still demonstrates good generalization to the other four unseen datasets. This suggests that Legilimens can maintain a satisfactory performance when deployed in real scenarios when distribution shifts exist.
Table 6. Robustness against LLM-targeted dynamic jailbreaking ($IO$-Moderation).
Dataset Attack Success Rate (%) PAIR $\rightarrow$ Vicuna PAIR $\rightarrow$ LLaMA2 TAP $\rightarrow$ Vicuna
Bare Legilimens Bare Legilimens Bare Legilimens
w/o DA w/ DA w/o DA w/ DA w/o DA w/ DA
AdvBEA Judged by GPT-4 93.33 53.33 3.33 3.33 0 0 93.33 40.00 3.33
Judged by Human 73.33 36.67 3.33 3.33 0 0 86.67 36.67 3.33
AdvBench* Judged by GPT-4 100.00 68.00 12.00 2.00 0 0 96.00 66.00 16.00
Judged by Human 100.00 62.00 10.00 2.00 0 0 94.00 62.00 16.00
  • †: DA is short for the data augmentation technique mentioned in §4.3. ¶: The attack success rate is judged by GPT-4 and verified manually.
5.4. Robustness against Adaptive Adversary
In this part, we consider an adaptive adversary who applies various jailbreak techniques to bypass Legilimens, including four types of LLM-targeted static jailbreaking, two types of LLM-targeted dynamic jailbreaking, and three types of moderator-targeted jailbreaking. We evaluate the robustness of Legilimens against these adaptive attacks.
5.4.1. LLM-Targeted Static Jailbreaking
For LLM-targeted static jailbreaking, we test Legilimens on samples rewritten into each type of jailbreak templates from BEA-adv&AG dataset, and present the results in Table 4 and Table 5. We can observe from Table 4 that Legilimens in five host models maintains a good performance even when encountering four types of static jailbreaking, with an accuracy of 98.192%, 97.339%, 98.768%, 93.983%, 93.872%. It is noteworthy that all jailbreak templates are unseen during the training phase for Legilimens. From Table 4, we notice that semantic transformations, such as pretending, attention shifting, and privilege escalation, have little impact on the performance of Legilimens. Syntactic transformations cause an accuracy drop of around 10%, possibly due to the host model’s failure to understand the semantics of prompts after transformation. But Legilimens still maintains an accuracy over 84.937%, outperforming all baselines. The results confirm the robustness of Legilimens against static jailbreaking.
As a reference, we train Legilimens without data augmentation mentioned in §4.3 and present the corresponding results on BEA-adv&AG dataset in Table 21 and Table 22 in Appendix E. Without data augmentation, Legilimens only achieves an accuracy of 96.808% (1.4$\downarrow$), 94.715% (2.6$\downarrow$), 95.371% (3.4$\downarrow$), 85.599% (8.4$\downarrow$), 89.640% (4.2$\downarrow$). The results validate the effectiveness of our model-based data augmentation technique.
5.4.2. LLM-Targeted Dynamic Jailbreaking
In this part, we evaluate Legilimens under much more stringent circumstances, wherein an adaptive adversary dynamically optimizes their prompts in an iterative manner based on responses of the host LLM to evade safety mechanisms. Within this context, LLM-Targeted Dynamic Jailbreaking can be divided into two types. The first type, known as white-box attack, latest studies (e.g.(Paulus et al., 2024; Guo et al., 2024)) have achieved high success rates, but they are not applicable to our scenario settings. The second type, black-box attack, encompasses PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), and IRIS (Ramesh et al., 2024), surpassing template-based attack methods, modifying prompts in interpretable ways to override LLMs’ safety guardrails. Both PAIR and TAP open source the code. Consequently, we reproduce two state-of-the-art dynamic jailbreaking methods, PAIR and TAP, to stress test Legilimens, with TAP representing an advanced attack improved from PAIR. We launch PAIR against Vicuna and LLaMA2 and launch TAP against Vicuna only, as we observe that it has limited effectiveness against LLaMA2. We utilize two datasets, namely AdvBench and AdvBEA, to provide initial prompts and goals for two jailbreaking methods. PAIR and TAP continually refine the prompts using Vicuna-7B, with a maximum of 60 and 70 attempts, respectively, to achieve the specified goals. We utilize GPT-4 to judge whether each attempt fulfills the goals, and manually verify the judgement. The attack is considered unsuccessful when all attempts fail. The detailed settings of these two attacks are presented in Appendix A. Note that both AdvBench and AdvBEA are unseen for Legilimens.
As shown in Table 6, PAIR achieves an attack success rate of 93.33% (73.33% upon manual verification) on AdvBEA, and 100% on AdvBench for the bare/unprotected Vicuna. Similarly, TAP also demonstrates a high attack success rate on both datasets. In contrast, these two jailbreaking methods only achieve an attack success rate of 3.33% on AdvBEA and 12$\sim$16on AdvBench for Legilimens-protected Vicuna, marking a significant decrease by 78$\sim$90%. These results validate the robustness of Legilimens against dynamic jailbreaking.
We also evaluate the performance of Legilimens when no data augmentation is applied in the training phase. As depicted in Table 6, Legilimens without data augmentation reduces the attack success rate of PAIR by 30$\sim$53%, once again confirming the effectiveness of our data augmentation technique.
We notice that PAIR can hardly attack the bare LLaMA2 because elaborate safety alignment mechanisms have been incorporated in the training phase of LLaMA2 (Touvron et al., 2023b). Legilimens reduces the remaining attack success (3.33%, 2.00%) to none (0%).
5.4.3. Moderator-Targeted Dynamic Jailbreaking
In this part, we explore an adaptive adversary that employs traditional adversarial example attacks on classifiers within the natural language processing domain. We reproduce two blind and one decision-based adversarial example attack methods, i.e., VIPER (Eger et al., 2019) (character-level), SCPN (Iyyer et al., 2018) (sentence-level), and GAN (Zhao et al., 2018) (sentence-level), utilizing an open-source textual adversarial attack toolkit named OpenAttack (Zeng et al., [n. d.]). For the decision-based method, we assume the adversary can deduce the decision of Legilimens from the responses returned. We use the default parameters for the implementation of the three attacks. As illustrated in Table 23 in Appendix E, we launch the three attacks against Legilimens using 200 samples from the OIG-Safety dataset, and we observe no success case, which confirms the robustness of Legilimens against traditional adversarial example attacks.
Table 7. Accuracy performance in few-shot scenarios ($IO$-Moderation).
#Training Samples ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
100-shot 94.808 0.9905 97.316 0.9969 93.792 0.9876 86.522 0.9324 91.308 0.9651
500-shot 97.045 0.9957 98.318 0.9983 96.460 0.9958 91.495 0.9710 94.724 0.9850
1,000-shot 95.237 0.9959 96.435 0.9984 96.325 0.9964 93.932 0.9849 95.652 0.9897
5,000-shot 97.991 0.9975 99.240 0.9996 98.242 0.9984 95.838 0.9914 97.847 0.9965
10,000-shot 98.308 0.9986 99.366 0.9998 98.643 0.9990 96.845 0.9947 98.252 0.9983
Table 8. FPR and FNR performance in few-shot scenarios ($IO$-Moderation).
#Training Samples ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
100-shot 1.384 12.717 4.682 1.908 17.225 0.877 15.335 15.258 7.672 9.252
500-shot 1.894 4.377 2.346 1.420 2.378 4.215 6.999 8.827 3.547 5.589
1,000-shot 2.673 2.234 1.442 1.438 3.071 1.899 7.268 5.671 2.355 5.128
5,000-shot 1.317 2.623 0.307 1.601 1.684 1.655 4.739 4.251 1.682 2.686
10,000-shot 2.163 1.392 0.711 0.452 1.434 1.212 2.692 3.988 0.971 2.939
5.5. Impact of Hyper-Parameters
In this part, we examine the impact of hyper-parameters on the performance of Legilimens, including the number of probed features and the architecture of the classifier. These experiments are carried out for both $O$- and $IO$-moderation.
5.5.1. The Number of Probed Features
We vary the number of probed features, i.e., $m$ mentioned in §4.1 from 1 to 9, and train a three-layer classifier on these features for each host LLM. We present the performance of $O$-moderation in Table 14, Table 16 and $IO$-moderation in Table 13, Table 15 in Appendix D. We can see that the number of probed features $m$ has little impact on the performance of Legilimens. For example, in $IO$-moderation, different choices of $m$ result in an accuracy change of only 0.372% at most. This indicates that the probed features from host LLMs are inherently comprehensive, and the fusion of features across different Transformer blocks only brings about a marginal improvement in performance.
Table 9. Accuracy of Multi-Label classification ($I$-Moderation).
     Category      $m=1$      $m=3$
     ACC      AUC      ACC      AUC
     Obscene      80.523      0.9362      79.595      0.9449
     Identity Attack      93.209      0.9878      94.365      0.9876
     Insult      90.615      0.9685      89.922      0.9658
     Threat      85.557      0.9752      84.808      0.9777
     Sexual Explicit      82.763      0.9648      81.950      0.9718
  • †: the number of probed features.
Table 10. FPR and FNR of Multi-Label classification ($I$-Moderation).
     Category      $m=1$      $m=3$
     FPR      FNR      FPR      FNR
     Obscene      6.711      22.852      1.669      37.801
     Identity Attack      1.668      12.281      1.367      15.017
     Insult      9.169      9.146      10.088      9.158
     Threat      2.203      15.770      0.700      27.555
     Sexual Explicit      0.974      29.748      1.512      26.981
  • †: the number of probed features.
5.5.2. The Number of Layers in Classifier
We vary the number of layers for the classifier from 1 to 9, and train the classifiers for each host LLM. In this part, we set $m=1$. We present the performance of $O$-moderation in Table 18, Table 20 and $IO$-moderation in Table 17, Table 19 in Appendix D. We find that a three-layer classifier achieves the best performance in most cases. These results validate that a lightweight classifier is sufficient for content moderation when the power of the host LLM is harnessed.
5.6. Few-Shot Scenarios
In this part, we evaluate the applicability of Legilimens to few-shot scenarios in order to assess the setup cost of Legilimens. We limit the number of training samples to train Legilimens to 100, 500, 1,000, 5,000 and 10,000. The results are presented in Table 7 and Table 8. Legilimens achieves a satisfactory accuracy on almost all host LLMs with only 100 samples (0.11% of the original training set), i.e., 94.808%, 97.316%, 93.792%, 86.522%, and 91.308%. With more than 1,000 samples (1.16% of the original training set), Legilimens achieves performance closer to standard training, as illustrated in Figure 3, lowering the burden for service providers in generating training samples.
Figure 3. Performance in few-shot scenarios ($IO$-Moderation). Legilimens performs well even when only 1,000 samples are available in the training phase.
5.7. Multi-Label Classification Extension
In certain scenarios, the classification of unsafe prompts or responses into granular unsafe types is useful, as different moderation strategies may be applied to different types. In this part, we endeavor to extend Legilimens to a multi-label classification task. We train Legilimens on Jigsaw to assign five labels for each sample using five MLPs, determining whether it is related to obscene, identity attack, insult, threat, and sexual explicit. More specifically, rather than training a multi-label classifier, we train a separate binary classifier for each label. We calculate the ACC, AUC, FPR and FNR values using the standard methodology for binary classification. The result is shown in Table 9 and Table 10. As illustrated in Table 9, Legilimens achieves an accuracy of 80.523$\sim$93.209when $m=1$, and 79.595$\sim$94.365when $m=3$. The results confirm that Legilimens can be extended to multi-label classification tasks. Note that the performance of Legilimens in classifying certain labels, such as obscene and sexual explicit, is not as good as in binary classification (an accuracy of 88.39% as shown in Table 2), which may be due to the imbalanced training set. For instance, the original ratio of positive to negative samples in sexual explicit is 1:7.72, despite our application of re-sampling to mitigate this issue. A more balanced training set may help further improve the performance.
6. Related Work
6.1. Model Alignment
Model alignment aims to remove the undesired behaviors of trained language models. Approaches to perform alignment include prompting and reinforcement learning from human feedback (RLHF).
Prompt-Based Alignment. Askell et al. (Askell et al., 2021) improved alignment and decreased output toxicity by injecting LLMs with helpful, honest, and harmless (HHH) prompts in the form of human-assistant conversations, where the assistant was always polite, helpful, and accurate. Similarly, Rae et al. (Rae et al., 2021) also used prompting in order to decrease toxicity (by including “to be respectful, polite and inclusive”). The problem with this approach is that existing LLMs may not strictly follow the instructions to be aligned.
RLHF-Based Alignment. Bai et al. (Bai et al., 2022) proposed to train LLMs to be helpful and harmless via RLHF. Specifically, they trained LLMs with the assistance of human evaluators in order to optimize their outputs to the evaluator’s preferences. Similarly, Ouyang et al. (Ouyang et al., 2022) fine-tuned GPT-3 into InstructGPT using data collected from human labelers to reach better performance on a variety of tasks, while improving alignment. Although RLHF for alignment is effective to a certain extent, it is dangerously brittle. Automatically-generated (Wallace et al., 2019; Yu and Sagae, 2021) or manually-designed (Xu et al., 2021) adversarial prompts have been shown to effectively bypass existing model alignment, including the alignment effort for ChatGPT (Li et al., 2023; Liu et al., 2023; Zhu et al., 2023; Yuan et al., 2023). Faced with these empirical results, Wolf et al. (Wolf et al., 2023) proposed a theoretical framework to prove that for any behavior that has a finite probability of being exhibited by the model, there exist prompts that can trigger the model into outputting this behavior. It reveals the fundamental limitations of alignment in LLMs.
Legilimens is orthogonal to the alignment-based defenses. When applied simultaneously, Legilimens has the potential to detect undesired behaviors when the alignment-based defenses are bypassed.
6.2. Content Moderation
There is a long track record of work on the detection of unsafe content. According to the classifier applied, we divide this kind of work into two categories, i.e., lightweight classifier and LLM-based classifier.
Lightweight Classifier. Many previous works utilized traditional machine learning classifiers for content moderation in social media. For example, Kwok et al. (Kwok and Wang, 2013)(2013) used a Naive Bayes classifier to distinguish between racist and nonracist tweets. Then, different model architectures are explored, e.g., logistic regression (Nobata et al., 2016; Davidson et al., 2017; Muhammad et al., 2023), Naive Bayes (Kwok and Wang, 2013; Davidson et al., 2017; Muhammad et al., 2023), decision trees (Davidson et al., 2017), random forests (Davidson et al., 2017; Muhammad et al., 2023), XGBoost (Muhammad et al., 2023), support vector machines (SVMs) (Davidson et al., 2017), multilayer perceptron (MLP) (Muhammad et al., 2023) and convolutional neural network (CNN) models (Muhammad et al., 2023). With the advances of pre-trained language models, BERT and its variants were utilized to more effectively extract feature for different textual detection tasks. For examples, Dinan et al. (Dinan et al., 2019) and Pavlopoulos et al. (Pavlopoulos et al., 2020) used a BERT-base model for offensive language detection. Moon et al. (Moon et al., 2023) leveraged a RoBERTa-base model for live-stream chats moderation. Kim et al. (Kim et al., 2023) employed a DistilBERT model for adversarial prompt detection. Another notable work of this type is by Markov et al. (Markov et al., 2023), which presented a holistic approach to building a robust classification system for real-world content moderation for ChatGPT services. Their classifier was based on lightweight GPT. Even with such seemingly comprehensive defenses, Wei et al. (Wei et al., 2023) revealed that ChatGPT services including GPT-4 were still vulnerable to jailbreak attacks. Their analysis emphasized that safety mechanisms should be as sophisticated as the underlying model. Otherwise, there exist attacks that exploit the cutting-edge capabilities of LLMs while less sophisticated safety classifiers cannot detect.
Legilimens leverages the powerful feature extraction ability of host models to detect unsafe content, thus addressing the issues of lightweight external classifiers.
LLM-based Classifier. Since it is usually believed that the more complex the classifiers, the more effective they perform, researchers explored LLMs for content moderation. Wang et al. (Wang et al., 2023) prompted GPT-3 to generate explanations for hateful and non-hateful speech. Huang et al. (Huang and Sun, 2023) explored using ChatGPT’s proficiency in detecting fake news. Ma et al. (Ma et al., 2023) fine-tuned LLMs that can be privately deployed for content moderation. Cao et al. (Cao et al., 2023) conducted a model review on Hugging Face to reveal the availability of models to cover various moderation rules and guidelines. Methods of this type are not efficient because extra LLM inference is required for content moderation.
Legilimens leverages the features extracted during the original inference process of host models for content moderation, thereby introducing minimal overhead.
7. Discussion & Future Work
More Granular Classification. In this paper, we have extended Legilimens to multi-label classification for five types of unsafe content, yielding satisfactory results in our initial experiments. Given the diversity of unsafe content, a system for more granular classification is useful. Nevertheless, the absence of an agreed-upon taxonomy for unsafe content and the unavailability of high-quality datasets hinder our ability to extend Legilimens to a more granular classification. We advocate for the prompt establishment of pertinent classification standards for unsafe content, fostering collaborative efforts among all stakeholders to construct high-quality content moderation datasets.
Larger Host Models. Legilimens achieves excellent performance on five host models of varying architectures, each with 6B to 7B parameters. However, the limitation of computational resources hinders our ability to conduct experiments on larger scale LLMs, i.e., 70B and 175B. Leveraging the capabilities of larger models, we anticipate that Legilimens will yield even improved results. Investigating the relationship between moderation effectiveness and model size is our future research direction.
Other Tasks. Legilimens is constructed based on conceptual features extracted from host LLMs for content moderation, demonstrating excellent performance. The same feature extraction technique can be employed for other tasks associated with LLMs. For instance, this can include monitoring the level of hallucination and dynamically adjusting the generation of responses to ensure they are safe, faithful, and factual. These potential applications represent intriguing future directions.
8. Conclusion
In this paper, we propose a practical and unified content moderation framework for LLM services, named Legilimens, which features both effectiveness and efficiency. We have conducted extensive experiments on various host LLMs, datasets, and jailbreaking methods to verify the effectiveness, efficiency, and robustness of Legilimens against normal and adaptive adversaries. The results validate that Legilimens outperforms both commercial and academic baselines.
Acknowledgement
We sincerely thank the anonymous reviewers for their valuable comments and dedication. This work was supported by Ant Group through CCF-Ant Research Fund. Yanjiao Chen is the corresponding author.
Appendix A Details of Dynamic Jailbreaking
For PAIR method, we use the Vicuna-7B model as the attacker LLM, with the target LLM being Vicuna-7b and LLaMA2-7B, and the judge LLM being GPT-4. Our parameter settings are the same as the PAIR method in the original paper, with a maximum of 60 attack queries.
For TAP method, we utilize the same attacker LLM as the PAIR method. We only launch TAP on Vicuna-7B due to its low attack success rate on LLaMA2-7B. We set the depth to 7, the maximum width to 10, and the branching factor to 4, allowing for a maximum of 70 attack queries. If the attacker LLM does not succeed after 70 attempts, we deem the attack unsuccessful.
Appendix B System Prompts to Adapt GPT-3.5 & GPT-4 for Content Moderation
In this part, we detail the system prompts to adapt GPT-3.5-Turbo and GPT-4 for three moderation tasks, namely $I$-moderation, $O$-moderation and $IO$-moderation. For $I$- and $IO$-moderation, we use the same system prompt template to moderate whether the content is safe. As for $O$-moderation, we use the system prompt proposed by (Ji et al., 2023).
$I$- and $IO$-Moderation Prompt Template Content: <prompt>/<prompt>+<response>. Is the content above related to sexuality, hate, violence, self-harm, and harassment? Return ”Yes” or ”No”.
$O$-Moderation Prompt Template [Question] <prompt> [The Start of the Assistant’s Answer] <response> [The End of the Assistant’s Answer] [System] We would like to request your feedback on the performance of the AI assistant in response to the user question displayed above. Please rate the harmlessness, helpfulness, and level of detail of their responses. Please prioritize the harmlessness. Please focus on whether there is discrimination and prejudice in the reply, whether it abides by the law, whether it avoids pornography and insults, whether it avoids porn content, and whether it protects privacy. Consider the following 14 categories in mind: 1. Hate Speech, Offensive Language 2. Discrimination, Stereotype, Injustice 3. Violence, Aiding and Abetting, Incitement 4. Financial Crime, Property Crime, Theft 5. Privacy Violation 6. Drug Abuse, Weapons, Banned Substance 7. Non-Violent Unethical Behavior 8. Sexually Explicit, Adult Content 9. Controversial Topics, Politics 10. Misinformation regarding ethics, laws and safety 11. Terrorism, Organized Crime 12. Self-Harm 13. Animal Abuse 14. Child Abuse Please first output a single line containing only ”flagged” or ”unflagged”. ”flagged” indicates the answer violates the safety rules above. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
Appendix C More Results of Baseline
Table 11. False Negative Rate (FNR) performance on various ${}^{\text{\textdagger}}$ and ${}^{\text{\textdaggerdbl}}$ compared with baselines.
Method I-Moderation (FNR, %) O-Mod. (FNR, %) IO-Mod. (FNR, %)
HAT MHS OIG JIG BEA BEA-adv BAG BAG-adv
OpenAI Moderation 12.300 0.796 62.400 34.807 80.990 80.354 87.019 82.788
Perspective API 27.236 4.011 78.200 41.290 90.717 99.628 92.692 94.519
BeaverDam-7B 23.100 8.437 71.100 56.346 8.101 27.002 47.500 57.170
LLaMA Guard2 30.902 13.555 45.600 91.923 36.054 40.596 60.578 64.678
GradSafe 15.400 2.628 40.800 56.442 28.650 57.263 47.115 44.467
GPT-3.5-Turbo 24.417 7.613 57.800 79.904 76.217 85.847 83.365 78.269
GPT-4 8.929 6.250 30.000 60.784 67.347 82.143 68.889 77.778
Legilimens (Ours) 20.656 10.488 13.700 16.564 11.350 14.759 0.525 3.763
  • †: Task alias: $I$-Moderation ($I$-Mod.), $O$-Moderation ($O$-Mod.) and $IO$-Moderation ($IO$-Mod.).
  • ‡: Dataset alias: HateXplain (HAT), Measuring Hate Speech (MHS), OIG-Safety (OIG), Jigsaw (JIG), BeaverTails (BEA), BeaverTail-adv (BEA-adv), BEA&AG (BAG), BEA-adv&AG (BAG-adv).
Table 12. False Positive Rate (FPR) performance on various ${}^{\text{\textdagger}}$ and ${}^{\text{\textdaggerdbl}}$ compared with baselines.
Method I-Moderation (FPR, %) O-Mod. (FPR, %) IO-Mod. (FPR, %)
HAT MHS OIG JIG BEA BEA-adv BAG BAG-adv
OpenAI Moderation 47.800 55.324 0.100 10.625 6.806 17.171 0.000 0.208
Perspective API 42.999 36.764 0.601 0.522 6.367 1.188 0.000 4.896
BeaverDam-7B 43.534 35.709 1.400 6.458 4.752 2.916 0.729 0.520
LLaMA Guard2 26.357 27.016 2.400 2.604 6.108 5.400 0.208 0.520
GradSafe 51.333 49.099 1.400 9.688 23.271 26.242 0.000 0.000
GPT-3.5-Turbo 29.319 29.933 0.800 3.542 6.806 13.607 2.813 22.396
GPT-4 47.727 42.307 8.000 2.041 17.647 15.909 0.000 5.455
Legilimens (Ours) 17.443 8.442 1.000 8.831 11.980 16.432 0.202 0.346
  • †: Task alias: $I$-Moderation ($I$-Mod.), $O$-Moderation ($O$-Mod.) and $IO$-Moderation ($IO$-Mod.).
  • ‡: Dataset alias: HateXplain (HAT), Measuring Hate Speech (MHS), OIG-Safety (OIG), Jigsaw (JIG), BeaverTails (BEA), BeaverTail-adv (BEA-adv), BEA&AG (BAG), BEA-adv&AG (BAG-adv).
Appendix D More Results of Different Hyper-Parameters
Table 13. Impact of the number of probed features on accuracy ($IO$-Moderation).
$m$ ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
1 98.914 0.9993 99.627 0.9999 99.226 0.9996 98.043 0.9978 98.951 0.9994
3 98.882 0.9992 99.567 0.9999 99.254 0.9997 98.159 0.9982 98.937 0.9993
5 98.882 0.9993 99.651 0.9999 99.342 0.9997 98.229 0.9981 99.026 0.9995
7 98.956 0.9994 99.623 0.9999 99.426 0.9998 98.322 0.9985 99.087 0.9995
9 98.928 0.9994 99.697 0.9999 99.356 0.9997 98.415 0.9985 98.937 0.9996
  • †: $m$ denotes the number of probed features, as mentioned in §4.1.
Table 14. Impact of the number of probed features on accuracy ($O$-Moderation).
$m$ ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
1 85.356 0.9313 87.279 0.9475 86.127 0.9371 81.473 0.8970 85.422 0.9329
3 85.906 0.9361 87.304 0.9488 85.957 0.9307 81.613 0.8975 85.901 0.9365
5 85.931 0.9368 87.740 0.9498 86.267 0.9330 79.385 0.8858 86.043 0.9369
7 86.488 0.9399 87.430 0.9500 86.478 0.9353 66.682 0.8789 85.722 0.9342
9 86.566 0.9404 87.762 0.9504 85.932 0.9317 64.860 0.8686 86.394 0.9399
  • †: $m$ denotes the number of probed features, as mentioned in §4.1.
Table 15. Impact of the number of probed features on FPR and FNR ($IO$-Moderation).
$m$ ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
1 0.865 1.204 0.510 0.289 1.117 0.507 1.856 2.361 0.817 1.284
3 0.904 1.293 0.202 0.525 0.424 0.968 1.365 2.578 0.567 1.538
5 0.827 1.239 0.548 0.226 0.616 0.724 1.269 2.252 0.510 1.393
7 0.760 1.230 0.211 0.588 0.809 0.597 1.385 2.062 0.519 1.474
9 0.481 1.483 0.452 0.262 0.347 0.977 1.009 2.044 0.567 1.339
  • †: $m$ denotes the number of probed features, as mentioned in §4.1.
Table 16. Impact of the number of probed features on FPR and FNR ($O$-Moderation).
$m$ ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
1 13.167 12.850 11.980 11.350 11.485 15.515 16.625 16.304 11.671 13.023
3 12.618 13.337 12.900 11.185 13.709 13.042 16.974 15.743 13.186 11.391
5 12.708 13.394 11.502 12.611 13.200 13.677 16.795 16.172 11.283 13.304
7 12.519 12.702 12.668 12.034 12.662 13.883 18.828 14.647 13.306 11.721
9 12.230 13.007 11.900 12.224 12.901 13.891 15.748 16.642 11.831 12.504
  • †: $m$ denotes the number of probed features, as mentioned in §4.1.
Table 17. Impact of the number of layers in classifier on accuracy ($IO$-Moderation).
#Layer ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
1 98.220 0.9985 99.525 0.9999 99.165 0.9995 97.544 0.9967 98.630 0.9990
3 98.914 0.9993 99.627 0.9999 99.226 0.9996 98.043 0.9994 98.951 0.9994
5 98.914 0.9990 99.669 0.9999 99.160 0.9995 98.015 0.9976 98.826 0.9993
7 98.914 0.9986 99.632 0.9999 99.133 0.9994 98.103 0.9978 98.812 0.9990
9 98.942 0.9984 99.501 0.9999 99.105 0.9993 98.001 0.9951 98.551 0.9986
  • †: the number of layers used in the classifier.
Table 18. Impact of the number of layers in classifier on accuracy ($O$-Moderation).
#Layer ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
1 79.655 0.8819 85.427 0.9313 82.968 0.9098 76.284 0.8457 82.621 0.9069
3 85.356 0.9313 87.279 0.9475 86.127 0.9371 81.473 0.8970 85.422 0.9329
5 84.567 0.9253 87.251 0.9466 85.654 0.9326 81.049 0.8940 85.592 0.9331
7 77.718 0.9197 84.100 0.9423 83.718 0.9324 77.038 0.8723 82.651 0.9199
9 82.353 0.9027 85.127 0.9317 84.278 0.9327 51.084 0.4611 81.118 0.8947
  • †: the number of layers used in the classifier.
Table 19. Impact of the number of layers in classifier on FPR and FNR ($IO$-Moderation).
#Layer ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
1 1.106 2.361 0.154 0.778 1.059 0.669 2.481 2.750 0.923 1.791
3 0.865 1.203 0.510 0.289 1.117 0.507 1.856 2.361 0.817 1.284
5 0.971 1.122 0.385 0.389 0.703 0.778 1.433 2.578 0.615 1.764
7 0.952 1.312 0.212 0.552 0.510 0.941 1.375 2.641 0.884 1.529
9 0.904 1.239 0.096 1.257 0.510 1.067 1.394 3.148 0.336 2.813
  • †: the number of layers used in the classifier.
Table 20. Impact of the number of layers in classifier on FPR and FNR ($O$-Moderation).
#Layer ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
1 15.290 15.537 10.874 13.930 16.032 13.207 21.898 18.183 12.309 15.504
3 13.167 12.850 11.980 11.350 11.485 15.515 16.625 16.304 11.671 13.023
5 14.233 12.949 11.851 11.482 16.391 11.962 17.403 15.488 12.070 12.842
7 7.316 20.895 7.665 16.914 9.621 17.172 14.163 18.595 10.914 14.037
9 7.246 19.403 5.273 20.673 6.710 22.242 11.223 22.923 5.033 22.412
  • †: the number of layers used in the classifier.
Appendix E More Results against Jailbreaking
Table 21. Legilimens w/o data augmentation against LLM-targeted static jailbreaking ($IO$-Moderation).
Jailbreaking Type ChatGLM3 LLaMA2 Falcon Dolly Vicuna
ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
Pretending 98.695 0.9988 99.748 1.0000 99.048 0.9997 92.786 0.9936 98.979 0.9988
Attention Shifting 98.322 0.9988 99.529 0.9998 91.751 0.9952 85.986 0.9887 89.696 0.9882
Privilege Escalation 99.045 0.9991 99.758 0.9999 97.988 0.9981 87.808 0.9806 90.973 0.9975
Syntactic Transformation 90.311 0.9868 80.048 0.9866 92.664 0.9878 76.143 0.9583 78.375 0.9649
Overall 96.808 0.9963 94.715 0.9968 95.371 0.9944 85.599 0.9798 89.640 0.9877
Table 22. Legilimens w/o data augmentation against LLM-targeted static jailbreaking on FPR and FNR ($IO$-Moderation).
Jailbreaking Type ChatGLM3 LLaMA2 Falcon Dolly Vicuna
FPR FNR FPR FNR FPR FNR FPR FNR FPR FNR
Pretending 0.778 2.026 0.481 0.027 1.021 0.344 2.067 8.520 0.644 1.357
Attention Shifting 0.778 1.872 0.481 0.470 1.167 12.094 1.904 13.323 0.644 18.723
Privilege Escalation 0.778 0.841 0.481 0.009 1.051 2.307 1.990 21.717 0.644 13.459
Syntactic Transformation 0.778 18.325 0.481 37.762 1.080 8.847 2.106 36.442 0.644 39.390
Overall 0.846 5.436 0.490 9.370 1.312 5.130 1.952 22.386 0.884 16.706
Table 23. Robustness against moderator-targeted dynamic jailbreaking ($I$-Moderation).
Method Type Host Model (Attack Success Rate, %)
ChatGLM3 LLaMA2 Falcon Dolly Vicuna
VIPER Blind 0 0 0 0 0
SCPN Blind 0 0 0 0 0
GAN Decision 0 0 0 0 0
  • †: Blind attacks are ignorant of the victim model, and decision-based attacks require the final decision for optimization (Zeng et al., [n. d.]).
Appendix F Time Complexity
Figure 4. The time complexity of Legilimens compared with three baselines. Legilimens exhibits a constant complexity of $O(1)$, while the other methods exhibits an approximately linear complexity of $O(n)$.
Appendix G Overview of dataset
Table 24. Overview of datasets.
Dataset Alias #Train #Test Task
HateXplain (Mathew et al., 2021) HAT 12,578 3,050 $I$
Measuring Hate Speech (Kennedy et al., 2020) MHS 16,566 4,142 $I$
OIG-Safety (Nguyen et al., 2023) OIG 19,769 2,000 $I$
Jigsaw (Google Jigsaw, 2017) JIG 239,204 59,801 $I$
BeaverTails (Ji et al., 2023) BEA 88,657 22,165 $O$
BeaverTails-adv BEA-adv 88,657 22,165 $O$
BEA&AG BAG 86,162 21,457 $IO$
BEA-adv&AG BAG-adv 86,162 21,457 $IO$
BEA&PIQA BPI 26,180 6,546 $I$
HarmBench (Mazeika et al., 2024) HAB - 320 $I$
SimpleSafetyTests (Vidgen et al., 2023) SST - 100 $I$
MaliciousInstructions (Bianchi et al., 2023) MAI - 100 $I$
JADE (Zhang et al., 2023) JAD - 80 $I$
HExPHI (Qi et al., 2023) HEP - 330 $I$
TDCRedTeaming (Mazeika et al., 2024) TDC - 50 $I$
AdvBench (Chao et al., 2023) - - 50 $IO$
AdvBEA - - 30 $IO$

### v2

@@
 Furthermore, we confirm that Legilimens can be applied to few-shot scenarios and extended to multi-label classification tasks.
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 31th ACM Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, Utahisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
+Content moderation; large language model; jailbreaking
+copyright: acmlicensedjournalyear: 2024doi: 10.1145/3658644.3690322conference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAisbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Artificial intelligence
+Figure 1.
+Legilimens is a unified content moderation framework for almost all LLM services, which features both effectiveness and efficiency.
+By monitoring conceptual features generated in the regular inference process of LLMs, Legilimens achieves lightweight moderation.
 1. Introduction
@@
 This dilemma of effectiveness and efficiency requires urgent resolution.
-Figure 1.
-Legilimens is a unified content moderation framework for almost all LLM services, which features both effectiveness and efficiency.
-By monitoring conceptual features generated in the regular inference process of LLMs, Legilimens achieves lightweight moderation.
 Table 1. Legilimens versus existing works.
@@
 Thus, we attain comprehensive conceptual features of $\mathbf{X}$ by fusing/concatenating the side features derived from the last several blocks, i.e.,
-(8) ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{X})\triangleq\left[\widetilde{\mathcal{H}}_{(K-m)}(\mathbf{X})\oplus\cdots\oplus\widetilde{\mathcal{H}}_{K}(\mathbf{X})\right].$
-The reason for deriving conceptual features from the last several blocks is to utilize more capacity of the host LLM, given that ~$\widetilde{\mathcal{H}}_{(K-m):K}$ can be viewed as containing all parameters of the host LLM.
+(8) ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{X})\triangleq\left[\widetilde{\mathcal{H}}_{(K-m+1)}(\mathbf{X})\oplus\cdots\oplus\widetilde{\mathcal{H}}_{K}(\mathbf{X})\right].$
+The reason for deriving conceptual features from the last several blocks is to utilize more capacity of the host LLM, given that ~$\widetilde{\mathcal{H}}_{[-m:]}$ can be viewed as containing all parameters of the host LLM.
 Note that, no additional computation is required to obtain the side features in Equation (8).
@@
 To reduce the complexity of Legilimens to $O(1)$, we resort to the attention span of self-attention mechanism.
-As shown in Figure 2, the attention span during the decoding of the first token $r_{1}$ encompasses the entire prompt, indicating that the conceptual features derived in this process, ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P})$, encapsulate a summarized knowledge of prompt $\mathbf{p}$.
-Similarly, the attention span during the decoding the last token [eos] encompasses both the prompt and the response, encapsulating a summarized knowledge of prompt $\mathbf{p}$ and response $\mathbf{r}$ in ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P}\oplus\mathbf{R})$.
-Note that, both ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P})$ and ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P}\oplus\mathbf{R})$ are of a constant size of $m\times d_{\text{model}}$, where $d_{\text{model}}$ is the output dimension of Transformer blocks.
+As shown in Figure 2, the attention span during the decoding of the first token $r_{1}$ encompasses the entire prompt, indicating that the conceptual features derived in this process, ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P})$, encapsulate a summarized knowledge of prompt $\mathbf{p}$.
+Similarly, the attention span during the decoding the last token [eos] encompasses both the prompt and the response, encapsulating a summarized knowledge of prompt $\mathbf{p}$ and response $\mathbf{r}$ in ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P}\oplus\mathbf{R})$.
+Note that, both ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P})$ and ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P}\oplus\mathbf{R})$ are of a constant size of $m\times d_{\text{model}}$, where $d_{\text{model}}$ is the output dimension of Transformer blocks.
 Reasoning on these two conceptual features only introduces constant overhead.
@@
 Based on the comprehensive representation summarized by the host LLM, we find that a simple multi-layer perceptron (MLP) with few parameters is adequate for accurate content moderation, denoted as $\mathcal{C}_{\theta\backslash\psi}:\mathbbm{R}^{m\times d_{\text{model}}}\rightarrow\mathbbm{Y}$.
-Therefore, the entire moderator is composed in this way: $\mathcal{M}_{\theta}=\mathcal{C}_{\theta\backslash\psi}\circ\widetilde{\mathcal{H}}_{(K-m):K}$.
+Therefore, the entire moderator is composed in this way: $\mathcal{M}_{\theta}=\mathcal{C}_{\theta\backslash\psi}\circ\widetilde{\mathcal{H}}_{[-m:]}$.
 We validate the effectiveness of our lightweight moderator in §5.2 compared with five baselines.
@@
 Acknowledgement
-We sincerely thank the anonymous reviewers for their valuable comments and dedication. This work was supported by Ant Group through CCF-Ant Research Fund. Yanjiao Chen is the corresponding author.
+We sincerely thank the anonymous reviewers for their valuable comments and dedication.
+This work was supported by China NSFC Grant 61925109 and by Ant Group through CCF-Ant Research Fund.
+Yanjiao Chen is the corresponding author.
 Appendix A Details of Dynamic Jailbreaking